# FlashKAT: Understanding and Addressing Performance Bottlenecks in the Kolmogorov-Arnold Transformer

### v2

@@ Abstract
The Kolmogorov-Arnold Network (KAN) has been gaining popularity as an alternative to the multi-layer perceptron (MLP) with its increased expressiveness and interpretability. Even so, the KAN suffers from being orders of magnitude slower due to its increased computational cost and training instability, limiting its applicability to larger-scale tasks. Recently, the Kolmogorov-Arnold Transformer (KAT) has been proposed, which can achieve FLOPs similar to the traditional Transformer with MLPs by leveraging Group-Rational KAN (GR-KAN). Unfortunately, despite the comparable FLOPs, our testing reveals that the KAT is still 123x slower in training speeds, indicating that there are other performance bottlenecks beyond FLOPs. In this paper, we conduct a series of experiments to understand the root cause of the slowdown in KAT. We uncover that the slowdown can be isolated to memory stalls, linked more specifically to inefficient gradient accumulations in the backward pass of GR-KAN. To address this memory bottleneck, we propose FlashKAT, which minimizes accesses to slow memory and the usage of atomic adds through a restructured kernel. Evaluations demonstrate that FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT, while reducing rounding errors in the computation of the gradients.
@@ 1 Introduction
The Kolmogorov-Arnold Network (KAN) has emerged as a popular alternative to the multi-layer perceptron (MLP) (liu2024kan). Rather than learning a fixed weight on each edge, the KAN instead learns a flexible nonlinearity. In doing so, the KAN is capable of increased expressivity and interpretability compared to the MLP. Initial works have primarily demonstrated the impressive capabilities of the KAN in scientific applications and equation modeling tasks (li2025u; coffman2025matrixkan; liu2024kan; koenig2024kan; wang2024kolmogorov). However, even with the superior performance of the KAN for these tasks, it has fallen short on larger-scale tasks in NLP and computer vision due to substantially increased parameters and computational cost, lack of GPU hardware optimization (e.g., for basis functions), and training instability (yu2024kan; le2024exploring; VisionKAN2024). For example, while computing an edge for an MLP only requires 2 FLOPS (1 for multiply and 1 for add), computing an edge for KAN may require up to 204 FLOPs, due to the nonlinearity (yang2024kolmogorov). Such a shortcoming necessitates significant research toward reducing the orders of magnitude performance gap between KAN and MLP.
@@
Recently, the Kolmogorov-Arnold Transformer (KAT) has managed to overcome some of the obstacles preventing the KAN from being deployed to large-scale tasks (yang2024kolmogorov). It achieves this by replacing each MLP in the Transformer with a Group-Rational KAN (GR-KAN), which augments the input to each linear layer with the safe Pade Activation Unit (PAU)(molina2019pad) in a grouped fashion (more details in Section 2). In doing so, each GR-KAN would learn a weighted summation of learned nonlinear functions. Unlike the traditional KAN, the GR-KAN is more computationally efficient, reducing the number of FLOPs to only slightly more than that of an MLP, and is constructed to better map to GPUs. By modifying the Transformer with GR-KAN, the resulting KAT is capable of achieving impressive results on computer vision tasks.
@@
However, even with the comparable number of FLOPs (and similar expected speed) offered by the KAT, in our characterization, the KAT is significantly slower compared to its MLP counterpart. For instance, our experiments reveal that a KAT (using GR-KAN) is 123x slower during training than an identically sized Vision Transformer (ViT, using MLP) (dosovitskiy2020image) on ImageNet-1K (russakovsky2015imagenet). This realization motivated an in-depth investigation to comprehensively understand the root causes of KAT’s performance bottleneck. Unlike previous approaches focused primarily on FLOPs, our methodology re-examines the performance issue from a memory-centric perspective, a novel angle not previously explored (yang2024kolmogorov), which uncovers that the slowdown is primarily due to a memory bottleneck rather than excessive computation. Specifically, we identify that the inefficient gradient accumulation method for the PAU coefficients, involving atomic adds, is the key contributor to this bottleneck (molina2019pad). This insight not only challenges assumptions about computation as the main constraint in the KAT but also provides a new direction for optimizing memory access patterns in similar architectures.
@@
To address the memory bottleneck, we propose FlashKAT, a substantially faster KAT that solves the GR-KAN (and by extension the PAU) shortcomings by not only avoiding accumulating gradients with atomic adds but also the unnecessary accesses to high-bandwidth memory (HBM). The method restructures data accesses and allows us to leverage shared memory to perform a sum reduction for gradient accumulations. By implementing our method as a GPU kernel, the resulting FlashKAT can achieve a training speedup of 86.5x compared with the state-of-the-art KAT. Furthermore, we demonstrate that our new kernel can reduce rounding errors in the gradient computations of the PAU coefficients by nearly an order of magnitude, thus improving training stability.
@@
As our method is a direct optimization of the PAU, a type of learnable rational activation function, the techniques and analysis also apply to alternative learnable rational activation functions, a complementary area of research to the KAT (molina2019pad; delfosse2021recurrent; delfosse2020rationals). Despite the clear need for efficient implementations, researchers in both the KAT and rational activation communities had not identified a way to accelerate these activations. FlashKAT provides this essential optimization, representing a nontrivial breakthrough that not only advances state-of-the-art but also inspires future research in the field.
@@
    Conducting a detailed investigation to understand the computational inefficiencies of the KAT, revealing that the primary performance bottleneck arises not from FLOPs, but from excessive memory traffic caused by inefficient gradient accumulation.
@@ 2 Background and Related Works
Recently, the KAN has been popularized as an alternative to the traditional MLP (liu2024kan). It is based on the Kolmogorov-Arnold Theorem (kolmogorov1961representation), which asserts that a multivariate continuous function on a bounded domain $f:[0,1]^{n}\rightarrow\mathbb{R}$ can be rewritten as a finite summation of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ as expressed in Equation 1:
@@
Although Equation 1 presents a theoretical form of the Kolmogorov-Arnold Theorem, the authors of (liu2024kan) demonstrated its effectiveness in a more generalized form by scaling the width and depth arbitrarily for a defined task. Suppose $n_{l}$ is the number of input nodes for layer $l$, where each input is represented by $x_{l,i}$. Each edge connection to layer $l+1$ has a learned activation $\phi_{l,j,i}(\cdot)$, where $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$. Then, we can represent each output node as
@@
Thus, if we have a $L$ layer KAN, the KAN can be represented as
@@
The most widely adopted method for representing $\phi_{l,j,i}(\cdot)$ has been with B-splines (liu2024kan), and although the B-spline-based KANs demonstrate promising results, they still have a host of shortcomings (yang2024kolmogorov; yu2024kan). These shortcomings consist of (1) an increased parameter and computational cost compared to the MLP alternative (e.g. increased FLOPs as summarized in Table 1), (2) B-splines not being GPU optimizable (e.g., requiring a recursive algorithm), and (3) training instability in larger networks (yang2024kolmogorov). Such shortcomings limit the KAN’s compatibility with the Transformer, restricting its application to larger-scale tasks like computer vision (vaswani2017attention).
@@
Unlike its predecessors, littered with shortcomings, the KAT succeeds in merging the KAN into a Transformer. It does so by developing a variation of the KAN called GR-KAN, which (1) shares coefficients across edge groups, reducing FLOPs, (2) leverages safe PAU activation functions, eliminating recursion, and (3) applies a variance-preserving initialization, increasing training stability (yang2024kolmogorov). Using the GR-KAN, the authors of (yang2024kolmogorov) construct the KAT by replacing the MLP with the GR-KAN, where each weight in the MLP is augmented with a nonlinear, learned activation function (i.e., a nonlinear edge in the KAN).
@@
First, rather than learning $d_{in}\cdot d_{out}$ activations, $\phi_{l,j,i}(\cdot)$, like a KAN, the GR-KAN groups inputs and only learns $\lfloor d_{in}/d_{g}\rfloor$ activations, $F(\cdot)$, where $d_{g}$ is the number of activations that share the same set of learned coefficients in a group. Each output $x_{l+1,j}$ is then a weighted summation of $F_{\lfloor i/d_{g}\rfloor}(x_{l,i})$ using learnable weights, $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$. The GR-KAN is expressed in Equation 5:
@@
Second, each activation in GR-KAN is a PAU (molina2019pad), which we will refer to as a group-wise rational function, composed of learnable coefficients $a_{i}$ and $b_{j}$ as follows:
@@
Third, a variance-preserving initialization is applied for stable learning (yang2024kolmogorov). The procedure for applying variance preserving weights requires (1) initializing the coefficients of $F(\cdot)$ to mimic a known activation function (ie. Swish, GeLU, Identity) and (2) initializing the weights in $\mathbf{W}$ according to $\mathcal{N}(0,\frac{\alpha}{d_{in}})$ whereby $\alpha=\frac{\mathbb{E}[\mathbf{F(x)}^{2}]}{Var[\mathbf{x}]}$ assuming $\mathbf{x}\sim\mathcal{N}(0,1)$.
Although the KAT is effective at vision tasks and is more computationally efficient than a ViT using the original KAN for training (a result of overcoming shortcomings (1), (2), and (3)), they still far short in terms of training speed compared to an identical ViT using an MLP, a shortcoming which has slowed further research. We demonstrate such results in the next section.
@@ 3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Insight 1: KAT is significantly slower than ViT during training.
@@
Although the KAT presents a promising alternative to the ViT, it still lags behind the ViT in training time. We illustrate this by comparing the combined forward and backward times for KAT-T (5.7M parameters), KAT-S (22.1M parameters), and KAT-B (86.6M parameters) against identically sized ViT-T, ViT-S, and ViT-B models, using the ImageNet-1k dataset for training (russakovsky2015imagenet). For the experiment, each KAT group-wise rational function contains 8 groups comprising of 6 numerator and 4 denominator coefficients, and all measurements are taken using an H200 GPU. The results, depicted in Figure 1, show the average time of 100 forward and backward passes, following 5 warmup passes. The data reveals that, for these forward/backward passes, KAT-T is 102x slower than ViT-T, KAT-S is 123x slower than ViT-S, and KAT-B is 116x slower than ViT-B.
Insight 2: FLOPs is not the performance bottleneck.
It is natural to assume that such a dropoff in training speed between the ViT and KAT results from the difference in FLOPs, as the KAT has more FLOPs than the ViT. However, by comparing the estimated FLOPs between the GR-KAN and MLP (Table 1), the only component in the ViT and KAT where they differ, we can dismiss such an assumption.
As can be seen, between the FLOPs expressions of the ViT and KAT, the main difference resides in the MLP activation requiring $\text{FuncFLOPs}\times d_{out}$ FLOPs and the GR-KAN requiring $(2m+2n+3)\times d_{in}$ FLOPs, both of which are negligible compared with the $2\times d_{in}\times d_{out}$ term. This suggests that the slightly increased FLOPs in the KAT may not be the source of the slowdown. To further verify this, we conduct an experiment that artificially increases the FLOPs inside the group-wise rational function in the KAT by placing an additional loop around the floating point computations. This would examine whether FLOPs bottleneck the performance. We measure performance metrics using Nvidia Nsight Compute [^1] The documentation for Nvidia Nsight Compute can be found at https://docs.nvidia.com/nsight-compute/ on an RTX 4060 Ti for the forward and backward pass. The dimension configuration is: input tensor, $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, numerator coefficients, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, denominator coefficients, $\mathbf{B}\in\mathbb{R}^{8\times 4}$, and the output gradient tensor, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$. We report these results in Table 2.
@@
From Table 2, we can immediately see that, even if the FLOPs scale to 8x of the original amount, the overall required ”Cycles” or ”Time” to complete the computation does not scale in either the forward or backward pass. Such results clearly demonstrate that the forward and backward pass kernels in KAT are not compute-bound, and we must look beyond FLOPs for performance bottlenecks.
Insight 3: Backward pass dominates execution time.
Results in Table 2 also demonstrate that the group-wise rational function dedicates a far greater amount of compute time to the backward pass than the forward pass. For instance, the forward pass takes 4.96 ms, whereas the backward pass requires 1.03 s, which is 207.7x as long. Therefore, to improve the training time of the KAT, performance issues for the backward pass must be identified and addressed. The backward pass can be broken down into three components: (1) loading values to compute the gradients, (2) computing the gradients, and (3) storing the gradients. Since we know the backward pass is not compute-bound from Insight 2, this means that (1) and (2) are to blame for the slowdown, both of which are memory issues.
Insight 4: Memory (and memory stall in particular) is the culprit.
@@
Figure 2: The warp‐state statistics for the KAT group-wise rational function backward pass. ”Computing - Selected” is where the warp does useful computation; all the others are memory stall (MS) states.
Focusing on the memory side for backward passes, we note that Table 2 reports very low L1, L2, and HBM throughput. At first glance, this is counterintuitive because a memory-bound process would usually mean a saturated memory bandwidth. However, further reasoning about the underlying hardware can provide insights. A GPU consists of a number of streaming multiprocessors (SMs). Typically, 32 threads are grouped in a warp and executed synchronously. Multiple warps can be scheduled for execution on an SM in a context-switching fashion. When one warp is waiting for a memory transfer, the SM can switch to doing computation for a different ”ready” warp (i.e., not waiting for memory), thus hiding the memory latency. However, suppose there are not enough ready warps. Then, the warp scheduler cannot entirely hide long latency or high memory contention, like with atomic operations that create sequential memory accesses. If an SM does not make adequate use of the memory hierarchy consisting of registers, shared memory, and global memory (a mapped memory composed of L1 cache, L2 cache, and HBM), the latency for each transfer becomes even more severe. For reference, on an A100, shared memory, L1, L2, and HBM have a latency of 29.0, 37.9, 261.5, and 466.3 clock cycles, respectively (luo2024benchmarking). Assuming that the group-wise rational function evaluation fails to hide latency, it makes sense that we would witness low memory utilization but still have a memory bottleneck. To validate this assumption, we examine the warp state statistics for the group-wise rational function backward pass using Nvidia Nsight Compute with the same configuration as before. These statistics provide the average cycles each warp stays in a given state for each instruction the kernel executes. We report the warp state statistics in Figure 2.
Figure 2 demonstrates that each warp spends the majority of its time waiting for various memory stalls to conclude. Each stall reported is related to a memory transfer, excluding the ”Selected” state, where the warp is selected to issue an instruction (i.e., computing). For instance, the ”Stall Long Scoreboard”, a stall on a memory transfer from global memory, accounts for 412x more time than the time spent in the ”Selected” state. As such, it becomes clear that the group-wise rational function slowdown is primarily due to memory stalls from memory transfers. This calls for a new method that can be aware of the memory hierarchy and avoid memory operations that create contention and restrict the latency hiding capabilities of the warp scheduler.

### v3

@@ Abstract
The Kolmogorov-Arnold Network (KAN) has been gaining popularity as an alternative to the multilayer perceptron (MLP) due to its greater expressiveness and interpretability. Even so, KAN suffers from training instability and being orders of magnitude slower due to its increased computational cost, limiting its applicability to large-scale tasks. Recently, the Kolmogorov-Arnold Transformer (KAT) has been proposed, achieving FLOPs comparable to traditional Transformer models with MLPs by leveraging Group-Rational KAN (GR-KAN). Unfortunately, despite the comparable FLOPs, our testing shows that KAT remains 123x slower during training, indicating that there are other performance bottlenecks beyond FLOPs. In this paper, we conduct a series of experiments to understand the root cause of the slowdown in KAT. We uncover that the slowdown can be isolated to memory stalls, linked more specifically to inefficient gradient accumulations in the backward pass of GR-KAN. To address this memory bottleneck, we propose FlashKAT, which minimizes accesses to slow memory and the usage of atomic adds through a restructured kernel. Evaluations show that FlashKAT achieves up to an 86.5x training speedup over state-of-the-art KAT while reducing rounding errors in gradient computation.
@@ 1 Introduction
The Kolmogorov-Arnold Network (KAN) has emerged as a popular alternative to the multilayer perceptron (MLP) (Liu et al. 2024). Rather than learning a fixed weight per edge, KAN instead learns a flexible nonlinearity. In doing so, KAN achieves greater expressivity and interpretability than MLP, which has led to impressive capabilities in scientific and equation modeling tasks (Li et al. 2025; Coffman and Chen 2025; Liu et al. 2024; Koenig et al. 2024; Wang et al. 2024). However, even with the superior performance of KAN for these tasks, it has fallen short on larger-scale tasks in NLP and computer vision due to substantially increased parameters and computational cost, lack of GPU hardware optimization (e.g., for basis functions), and training instability (Yu et al. 2024; Le et al. 2024; Chen et al. 2024). For example, while computing an edge for an MLP only requires 2 FLOPS (1 for multiply and 1 for add), computing an edge for a KAN may require up to 204 FLOPs, due to the nonlinearity (Yang and Wang 2024). This shortcoming necessitates significant research to reduce the orders of magnitude performance gap between KANs and MLPs.
@@
Recently, the Kolmogorov-Arnold Transformer (KAT) has overcome some of the obstacles that prevent KAN from being deployed to large-scale tasks (Yang and Wang 2024). It achieves this by replacing each MLP in the Transformer with a Group-Rational KAN (GR-KAN), which augments the input to each linear layer with the safe Pade Activation Unit (PAU) (Molina et al. 2019) in a grouped fashion (more details in Section 2). In doing so, each GR-KAN learns a weighted summation of learned nonlinear functions. Unlike KAN, GR-KAN is more computationally efficient, reducing FLOPs to only slightly more than MLPs, and is designed to better map to GPUs. By modifying the Transformer with GR-KAN, the resulting KAT achieves impressive results on computer vision tasks.
@@
However, even with the comparable number of FLOPs (and similar expected speed) offered by KAT, in our characterization, KAT is significantly slower than its MLP counterpart. For instance, our experiments reveal that a KAT (using GR-KAN) is 123x slower during training than an identically sized Vision Transformer (ViT, using MLP) (Dosovitskiy et al. 2020) on ImageNet-1K (Russakovsky et al. 2015). This realization motivated an in-depth investigation to understand the KAT’s performance bottleneck. Unlike previous approaches focused primarily on FLOPs, we re-examine the performance issue from a memory-centric perspective, a novel angle not previously explored (Yang and Wang 2024), which uncovers that the slowdown is primarily due to a memory bottleneck rather than excessive computation. Specifically, we identify the inefficient gradient accumulation method for the PAU coefficients, involving atomic adds, as the key contributor (Molina et al. 2019). This insight not only challenges assumptions that computation is the primary constraint in KAT but also offers a new direction for optimizing memory access patterns in similar architectures.
@@
To address the memory bottleneck, we propose FlashKAT, a substantially faster KAT that solves GR-KAN’s (and by extension PAU’s) shortcomings by not only avoiding accumulating gradients with atomic adds but also the unnecessary accesses to high-bandwidth memory (HBM). The method restructures data accesses and enables more effective use of shared memory for gradient accumulations. By implementing our method as a GPU kernel, FlashKAT achieves a training speedup of up to 86.5x compared with state-of-the-art KAT. Furthermore, we demonstrate that our new kernel reduces rounding errors in the PAU coefficient gradient computations by nearly an order of magnitude, thereby improving training stability.
@@
As our method directly optimizes PAU, a learnable rational activation, the techniques and analysis also apply to alternative learnable rational activations, a complementary area of research to KAT (Molina et al. 2019; Delfosse et al. 2020, 2021; Biswas et al. 2021). Despite the clear need for efficient implementations, researchers in both the KAT and rational activation communities had not identified a way to accelerate these activations. FlashKAT provides this essential optimization, representing a nontrivial breakthrough that not only advances the state of the art but also inspires future research in the field.
@@
    Conducting a detailed investigation to understand the computational inefficiencies of KAT, revealing that the primary performance bottleneck arises not from FLOPs, but from suboptimal memory accesses caused by inefficient gradient accumulation.
@@ 2 Background and Related Works
Recently, KAN has been popularized as an alternative to MLP (Liu et al. 2024). It is based on the Kolmogorov-Arnold Theorem (Kolmogorov 1961), which asserts that a multivariate continuous function on a bounded domain $f:[0,1]^{n}\rightarrow\mathbb{R}$ can be rewritten as a finite summation of univariate continuous functions, $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$ as expressed in Equation 1:
@@
Although Equation 1 presents a theoretical form of the Kolmogorov-Arnold Theorem, (Liu et al. 2024) demonstrated its effectiveness in a more generalized form by scaling the width and depth arbitrarily for a defined task. Suppose $n_{l}$ is the number of input nodes for layer $l$, where each input is represented by $x_{l,i}$. Each edge connection to layer $l+1$ has a learned activation $\phi_{l,j,i}(\cdot)$, where $i\in[1,n_{l}]$ and $j\in[1,n_{l+1}]$. Then, we can represent each output node as
@@
Thus, an $L$ layer KAN can be represented as
@@
The most widely adopted method for representing $\phi_{l,j,i}(\cdot)$ has been with B-splines (Liu et al. 2024), and although B-spline-based KANs demonstrate promising results, they still have a host of shortcomings (Yang and Wang 2024; Yu et al. 2024). These shortcomings consist of (1) increased parameter and computational cost compared to MLP (e.g., increased FLOPs as summarized in Table 1), (2) B-splines not being GPU optimizable (e.g., requiring a recursive algorithm), and (3) training instability in large networks (Yang and Wang 2024). Such shortcomings limit KAN’s compatibility with the Transformer (Vaswani et al. 2017), restricting its applicability to large-scale tasks like computer vision.
@@
Unlike its predecessors, littered with shortcomings, KAT succeeds in merging KAN into a Transformer. It does so by developing a new KAN variation called GR-KAN, which (1) shares coefficients across edge groups, reducing FLOPs, (2) leverages safe PAU activation functions, eliminating recursion, and (3) applies a variance-preserving initialization, increasing training stability (Yang and Wang 2024). In KAT, each MLP is replaced with a GR-KAN, causing each MLP weight to be augmented with a nonlinear learned activation function (i.e., a nonlinear edge in KAN).
@@
First, rather than learning $d_{in}\cdot d_{out}$ activations, $\phi_{l,j,i}(\cdot)$, like KAN, GR-KAN groups inputs and only learns $\lfloor d_{in}/d_{g}\rfloor$ activations, $F(\cdot)$, where $d_{g}$ is the number of activations that share the same set of learned coefficients in a group. Each output $x_{l+1,j}$ is then a weighted summation of $F_{\lfloor i/d_{g}\rfloor}(x_{l,i})$ using learnable weights, $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$. GR-KAN is expressed in Equation 5:
@@
Second, each GR-KAN activation is a safe PAU (Molina et al. 2019), which we will refer to as a group-wise rational function, composed of learnable coefficients $a_{i}$ and $b_{j}$ as follows:
@@
Third, a variance-preserving initialization is applied for stable learning (Yang and Wang 2024). The procedure for applying variance preserving weights requires (1) initializing the coefficients of $F(\cdot)$ to mimic a known activation function (ie. Swish) and (2) initializing the weights in $\mathbf{W}$ according to $\mathcal{N}(0,\frac{\alpha}{d_{in}})$ whereby $\alpha=\frac{\mathbb{E}[\mathbf{F(x)}^{2}]}{Var[\mathbf{x}]}$ assuming $\mathbf{x}\sim\mathcal{N}(0,1)$.
Although KAT is effective at vision tasks and is more computationally efficient than a version of ViT with KAN (a result of overcoming shortcomings (1), (2), and (3)), they still far short in terms of training speed compared to an identical ViT using MLP, a shortcoming which has slowed further research. We demonstrate such results in the next Section.
@@ 3 Beyond FLOPs: Understanding Performance Bottlenecks of KAT
Insight 1: KAT Is Significantly Slower Than ViT During Training
@@
Although KAT presents a promising alternative to ViT, it still lags behind ViT in training time. We illustrate this by comparing the combined forward and backward times for KAT-T (5.7M parameters), KAT-S (22.1M parameters), and KAT-B (86.6M parameters) against identically sized ViT-T, ViT-S, and ViT-B models, using the ImageNet-1k dataset for training (Russakovsky et al. 2015). For the experiment, each KAT group-wise rational function contains 8 groups comprising of 6 numerator and 4 denominator coefficients, and all measurements are taken using an H200 GPU. The results, depicted in Figure 1, show the average time of 100 forward and backward passes, following 5 warmup passes. The data reveals that, for these forward/backward passes, KAT-T is 102x slower than ViT-T, KAT-S is 123x slower than ViT-S, and KAT-B is 116x slower than ViT-B.
Insight 2: FLOPs Is Not the Performance Bottleneck
It is natural to assume that such a drop-off in training speed between ViT and KAT results from the difference in FLOPs, as KAT has more FLOPs than ViT. However, by comparing the estimated FLOPs for GR-KAN and MLP (Table 1), the only component in ViT and KAT where they differ, we can dismiss this assumption.
As can be seen, between the FLOPs expressions of ViT and KAT, the main difference resides in the MLP activation requiring $\text{FuncFLOPs}\times d_{out}$ FLOPs and the GR-KAN learnable activation requiring $(2m+2n+3)\times d_{in}$ FLOPs, both of which are negligible compared with the $2\times d_{in}\times d_{out}$ term. This suggests that the slight increase in FLOPs in KAT may not be the cause of the slowdown. To further verify this, we conduct an experiment that artificially increases the FLOPs inside the group-wise rational function of KAT by placing an additional loop around the floating-point computations. This would examine whether FLOPs bottleneck the performance. We measure performance metrics using Nvidia Nsight Compute [^1] The documentation for Nvidia Nsight Compute can be found at https://docs.nvidia.com/nsight-compute/ on an RTX 4060 Ti for the forward and backward pass. The dimension configuration is: input tensor, $\mathbf{X}\in\mathbb{R}^{1024\times 197\times 768}$, numerator coefficients, $\mathbf{A}\in\mathbb{R}^{8\times 6}$, denominator coefficients, $\mathbf{B}\in\mathbb{R}^{8\times 4}$, and the upstream gradient tensor, $\mathbf{dO}\in\mathbb{R}^{1024\times 197\times 768}$. We report these results in Table 2.
@@
From Table 2, we can immediately see that, even if the FLOPs scale to 8x of the original amount, the overall required “Cycles” or “Time” to complete the computation does not scale in either the forward or backward pass. Such results clearly demonstrate that the forward and backward pass kernels in KAT are not compute-bound, and we must look beyond FLOPs to identify performance bottlenecks.
Insight 3: The Backward Pass Dominates Execution Time
Results in Table 2 also demonstrate that the group-wise rational function dedicates a far greater amount of compute time to the backward pass than the forward pass. For instance, the forward pass takes 4.96 ms, whereas the backward pass requires 1.03 s, which is 207.7x as long. Therefore, to reduce KAT’s training time, performance issues in the backward pass must be identified and addressed. The backward pass can be broken down into three components: (1) loading values to compute the gradients, (2) computing the gradients, and (3) storing the gradients. Since we know the backward pass is not compute-bound from Insight 2, this means that (1) and (3) are to blame for the slowdown, both of which are memory issues.
Insight 4: Memory (and Memory Stall in Particular) Is the Culprit
@@
Figure 2: The warp‐state statistics for the KAT group-wise rational function backward pass. “Computing - Selected” is where the warp does useful computation; all the others are memory stall (MS) states.
Focusing on the memory side for backward passes, we note that Table 2 reports very low L1, L2, and HBM throughput. At first glance, this is counterintuitive because a memory-bound process would typically imply a saturated memory bandwidth. However, further reasoning about the underlying hardware can provide insights. A GPU consists of multiple streaming multiprocessors (SMs). Typically, 32 threads are grouped in a warp and executed synchronously. Multiple warps can be scheduled for execution on an SM in a context-switching fashion. When one warp is waiting for a memory transfer, the SM can switch to doing computation for a different “ready” warp (i.e., not waiting for memory), thereby hiding memory latency. However, suppose there are not enough ready warps. Then, the warp scheduler cannot entirely hide long latencies or high memory contention, like with atomic operations that create sequential memory accesses. If an SM does not make adequate use of the memory hierarchy consisting of registers, shared memory, and global memory (a mapped memory composed of L1 cache, L2 cache, and HBM), the latency for each transfer becomes even more severe. For reference, on an A100, shared memory, L1, L2, and HBM have latencies of 29.0, 37.9, 261.5, and 466.3 clock cycles, respectively (Luo et al. 2024). Assuming that the group-wise rational function evaluation fails to hide latency, it makes sense that we would witness low memory utilization yet still experience a memory bottleneck. To validate this assumption, we examine the warp state statistics for the group-wise rational function backward pass using Nvidia Nsight Compute with the same configuration as before. These statistics report the average number of cycles each warp remains in a given state for each instruction the kernel executes. We report the warp state statistics in Figure 2.
Figure 2 demonstrates that each warp spends the majority of its time waiting for various memory stalls to conclude. Each reported stall corresponds to a memory transfer, excluding the “Selected” state, in which the warp is selected to issue an instruction (i.e., compute). For instance, the “Stall Long Scoreboard”, a stall on a memory transfer from global memory, accounts for 412x the time spent in the “Selected” state. As such, it becomes clear that the group-wise rational function slowdown is primarily due to memory transfer stalls. This calls for a new method that is aware of the memory hierarchy and avoids memory operations that create contention, limiting the warp scheduler’s latency-hiding capabilities.